OpenStudy (anonymous):

The population of town A is double that of town B but it is decreasing at the rate of 5% per year. The population of town B is increasing at the rate of 5% per year. In how many years will the population of town B be double that of town A

OpenStudy (hba):

B0*(1.05)^n = 2A0*(.95)^n or B0*(1.05)^n = 2*2*B0*(.95)^n or (1.05/.95)^n = 4 or n = ln(4) / ln(1.05/.95) or n = 13.851383446 years

OpenStudy (hba):

Another Solution : Solve for t such that B(t) = 2 * A(t) B0 * 1.05^t = 2 * (A0 * 0.95^t) Substitute to eliminate A0: B0 * 1.05^t = 2 * (2 * B0 * 0.95^t) B0 * 1.05^t = 4 * B0 * 0.95^t 1.05^t = 4 * 0.95^t log(1.05^t) = log(4 * 0.95^t) log(1.05^t) = log(4) + log(0.95^t) t * log(1.05) = log(4) + t * log(0.95) t * log(1.05) - t * log(0.95) = log(4) t * (log(1.05) - log(0.95)) = log(4) t = log(4) / (log(1.05) - log(0.95))

OpenStudy (hba):

@stefffi

OpenStudy (anonymous):

@hba thanks a bunch :) especially for providing multiple solutions

OpenStudy (hba):

@stefffi You Are Welcome :)